OpenStudy (anonymous):

A bicycle wheel is 24" diameter rotates 215 degrees. What distance has the bike traveled?

OpenStudy (anonymous):

we can us the arc formula

OpenStudy (campbell_st):

the formula for the circumference is \[C = \pi \times d\] the fraction of the circle that the bike wheel turns is \[\frac{215}{360}\] find the circumference then multiply by the fraction

OpenStudy (anonymous):

how would I put it into the arc formula @mack7

OpenStudy (ribhu):

\[(215/360)*24\pi\]

OpenStudy (anonymous):

\[\frac{ \theta }{ 360 }\times2*\pi*r\]

OpenStudy (anonymous):

\[\theta=215\]

OpenStudy (ribhu):

here theta would be 215 degrees

OpenStudy (anonymous):

r=12

OpenStudy (anonymous):

put the value in the equation you will get the answer

OpenStudy (anonymous):

i think it helped you

OpenStudy (anonymous):

Can you help me solve it, Im not good with fractions

OpenStudy (anonymous):

these are the steps

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

please identify the best response to help others

OpenStudy (anonymous):

I'm supposed to get 45 in the end

OpenStudy (anonymous):

24 is in meter or not

OpenStudy (anonymous):

24 inches

OpenStudy (anonymous):

so first you will have to convert 24 inches to meters

OpenStudy (campbell_st):

you don't have to convert anything read the initial post... ot explains you need 1. to find the circumference of the wheel. 2. multiply the answer, the circumference, by 215/360 the reason is a full revolution is 360 degrees... the fraction it turns is 215/360 try it and see if you get your answer of 45...

OpenStudy (anonymous):

I figured it out